OpenStudy (anonymous):

z^2=i. Solve in a+bi form. I have been teaching myself the complex numbers material and I am not sure how to do this.

OpenStudy (anonymous):

there are two solutions to this. draw the unit circle in the complex plane. i sits right up top at an angle of \[\frac{\pi}{2}\] half that angle is \[\frac{\pi}{4}\] so find the coordinates on the unit circle with angle of \[\frac{\pi}{4}\] and that is your first answer

OpenStudy (anonymous):

second answer is directly across from the first answer on the unit circle

OpenStudy (anonymous):



OpenStudy (anonymous):

at an angle of \[\frac{\pi}{4}\] you get \[\frac{\sqrt{2}}{2}+\frac{\sqrt{2}}{2}i\] and that is one solution

OpenStudy (anonymous):

Hmm confusing. Is this using eulers formula?

OpenStudy (anonymous):

i guess

OpenStudy (anonymous):

the idea is this. when you square a complex number you square the absolute value and double the angle.

OpenStudy (anonymous):

so if you want the square root, you work backwards. now the absolute value of i is just 1, therefore you are working on the unit circle. so you can pretty much visualize the answer. half way between 1 and i on the unit circle is the point i drew

OpenStudy (anonymous):

Oh, okay. I was looking at notes and I didn't realize those brackets were still absolute values. Abs(z)=r=1

OpenStudy (anonymous):

right. and there is another one as well, the one right across the circle. since you are taking square roots there will be two answers.

OpenStudy (anonymous):

hm r will remain 1 but how did you get to a+bi

OpenStudy (anonymous):

i suppose you could say \[i=\cos(\frac{\pi}{2})+i\sin(\frac{\pi}{2})\] and then divide the angle by 2 to get your answer

OpenStudy (anonymous):

you have to know the point on the unit circle whose angle is \[\frac{\pi}{4}\]

OpenStudy (anonymous):

yeah got that

OpenStudy (anonymous):

there is no way to do this without the trig i don't think. and the other one is right across at \[-\frac{\sqrt{2}}{2}-\frac{\sqrt{2}}{2}i\]

OpenStudy (anonymous):

hah im sorry but still lost. I know that i=Cos(pi/2)+isin(pi/2)

OpenStudy (anonymous):

ok if you know that you are almost done. you want the two square roots, so divide the angle by 2

OpenStudy (anonymous):

right, but why do you just divide by two. Why wouldn't it be the sqrt(Pi/2)

OpenStudy (anonymous):

you get \[\cos(\frac{\pi}{4})+i\sin(\frac{\pi}{4})\]

OpenStudy (anonymous):

oh because when you multiply two complex numbers together you 1) multiply the modulus (absolute value) 2) ADD the angles

OpenStudy (anonymous):

and therefore if you want to raise a complex number to a power you 1) raise the absolute value to the power and 2) MULTIPLY the angles (not raise the angles to a power)

OpenStudy (anonymous):

and finally if you want to find the roots of a complex number you 1) take the root of the absolute value and 2) DIVIDE the angle

OpenStudy (anonymous):

gonna write that down! Well I think I almost got it. I am gonna think it over more. I REALLY appreciate your time and help.

OpenStudy (anonymous):

btw you can check by computing that \[(\frac{\sqrt{2}}{2}+\frac{\sqrt{2}}{2}i)^2=i\]

OpenStudy (anonymous):

might be worth the arithmetic to see how it works out.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

mathematica just spit out a simplified thing. So you mean do it by hand?

OpenStudy (anonymous):

yes i mean write \[(\frac{\sqrt{2}}{2}+\frac{\sqrt{2}}{2}i)(\frac{\sqrt{2}}{2}+\frac{\sqrt{2}}{2}i)\]

OpenStudy (anonymous):

multiply out and see what you get

OpenStudy (anonymous):

okay I will do that but couldn't you follow your rule above and do Abs(z1)(Abs(z2)=r*r and do Pi/4+Pi/4?

OpenStudy (anonymous):

well sure but that is how we got it to begin with.

OpenStudy (anonymous):

just verifies that you divided by two and took the square root of 1

OpenStudy (anonymous):

This stuff is crazy.

OpenStudy (anonymous):

I have to tackle ln(3+4i) next if you aren't sick of this yet.

OpenStudy (anonymous):

we wrote \[i=\cos(\frac{\pi}{2})+i\sin(\frac{\pi}{2})\] and took half the angle. so it stands to reason if you double the angle we took half of you will get back to what you started with. i meant multiply out in standard \[a+bi\] form to see that it actually works. here is a nice picture from wolfram solving \[z^6=1\] you can see that the solutions are evenly spaced about the unit circle http://www.wolframalpha.com/input/?i=z^6%3D1

OpenStudy (anonymous):

So the power determines the number of solutions?

OpenStudy (anonymous):

yes. fundamental theorem of algebra says a poly of degree n has n solution. if you want to solve \[z^n=1\] there will be n distinct solution, evenly spaced about the unit circle

OpenStudy (anonymous):

as for Log, it is \[\log(z) = \ln(|z|)+i \theta\]

OpenStudy (anonymous):

\[|3+4i|=5\] by pythagoras, the angle i don't know. so i guess you could write \[\log(3+4i)=\ln(5)+i\tan^{-1}(\frac{4}{3})\] or you could actually compute the angle with a calculator